A buffer is prepared by mixing 1.00 gram imidazole ( a weak base, B), with 2.00 gram imidiazole hydrochloride (conjugate weal acid BH+) and diluting to 122.0 mL. Calculate the pH of the solution if 1.8 mL of 1.18 M HClO4. are added to the solution. pKa = 6.993, FM of imidazole = 68.08, FM of imidazole hydrochloride = 104.54.
Answer #1

mole of imidazole = mass / molar mass = 1.00 g / 68.08 g / mole = 0.01469 mole.

mole of imidiazole hydrochloride = mass / molar mass = 2.00 g /  104.54 g / mole = 0.0191 mole.

1.8 mL of 1.18 M HClO4 = 0.0018 L * 1.18 mole / L = 2.124 * 10^-3 mole.

after addition of HClO4,

mole of imidazole = 0.01469 - 2.124 * 10^-3 = 0.0126 mole.

and

mole of imidiazole hydrochloride = (0.0191 + 2.124 * 10^-3) = 0.0212 mole.

pKb = 14 - pKA = 14 - 6.993 = 7.007

imidazole is weak base.

thus

Using henderson equation,

pOH = pKb + log [salt] / [base]

or

pOH = 7.007 + log (0.0212 / 0.0126) = 7.233

PH = 14 - 7.233 = 6.767

PH = 6.767
Answer #2

mole of imidazole = mass / molar mass = 1.00 g / 68.08 g / mole = 0.01469 mole.

mole of imidiazole hydrochloride = mass / molar mass = 2.00 g /  104.54 g / mole = 0.0191 mole.

1.8 mL of 1.18 M HClO4 = 0.0018 L * 1.18 mole / L = 2.124 * 10^-3 mole.

after addition of HClO4,

mole of imidazole = 0.01469 - 2.124 * 10^-3 = 0.0126 mole.

and

mole of imidiazole hydrochloride = (0.0191 + 2.124 * 10^-3) = 0.0212 mole.

pKb = 14 - pKA = 14 - 6.993 = 7.007

imidazole is weak base.

thus

Using henderson equation,

pOH = pKb + log [salt] / [base]

or

pOH = 7.007 + log (0.0212 / 0.0126) = 7.233

PH = 14 - 7.233 = 6.767

PH = 6.767